Consider the following Keynesian consumption model: ? = ?̅ + ??, where ?̅ = 200 and ? = 0.4. Calculate the amount of consumption (?) and the average propensity to consume (APC) when the level of income is 400.

a) C=160, APC=0.4
b) C=160, APC=0.9
c) C=360, APC=0.4

d) C=360, APC=0.9
Answer #1

Correct option - (D)

Given, autonomous consumption c' = 200

Income, Y = 400

MPC, c = 0.4

Consumption, C = 200 + (0.4) (400)

= 200 + 160 = 360

Consumption, C = 360

Now, APC = C/Y = 360/400 = 0.9

Therefore, APC = 0.9